OpenStudy (anonymous):

Solve for x: −5|x + 1| = 10 x = 0 x = −3 and x = 1 x = −1 and x = 3 No solution @DarkBlueChocobo

OpenStudy (nurali):

−5|x + 1| = 10 |x + 1| = 10/-5 |x + 1| = -2 x+1=-(-2) or x+1=-2 x+1=2 or x+1=-2 x=1 or x=-3

OpenStudy (darkbluechocobo):

^^ He is right :P